The magnitude of the velocity of a particle which moves along the curve x=2 sin ??¡3t,y=2 cos ??¡3t and z=8t at any time t>0 is ?
1
Expert's answer
2019-11-11T10:39:35-0500

Curve given is

"x=2sin(a^{3t})"

"y=2cos(a^{3t})"

"z=8t"

"v_x=dx/dt=2.cos(a^{3t}).a^{3t}\ln a.d(3t)/dt"

"v_x\u200b=2.cos(a^{3t}).a^{3t}.\\ln a.3"

"v_x=6.cos(a^ {3t} ).a^{3t}.\ln a" -------(i)

"v_y=dy/dt=-2.sin(a^{3t}).a^{3t}\ln a.d(3t)/dt"

"v_y=-2.sin(a^{3t}).a^{3t}.\ln a.3"

"v_y=-6.sin(a^{3t}).a^{3t}.\ln a" ---------(ii)

"v_z= dz\/dt= 8" ------(iii)

"v = v_x \hat{i}+v_y\hat{j}+v_z\hat{k}"

"|v| = \sqrt{v_x^2 + v_y^2+v_z^2}"

"|v|=\sqrt{(6.cos(a^ {3t} ).a^{3t}.\ln a)^2+(6.sin(a^ {3t} ).a^{3t}.\ln a)^2+ 8^2}"

"|v|=\sqrt{(6.a^{3t}.\ln a)^2(sin(a^{3t})^2+cos(a^{3t})^2)+8^2}"

"|v| =\sqrt{36.a^{6t}(\ln a)^2+64}" (Answer)